# VLN-MME : Diagnosing MLLMs as Language-guided Visual Navigation agents

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities across a wide range of vision-language tasks. However, their performance as embodied agents, which requires multi-round dialogue spatial reasoning and sequential action prediction, needs further exploration. Our work investigates this potential in the context of Vision-and-Language Navigation (VLN) by introducing a unified and extensible evaluation framework to probe MLLMs as zero-shot agents by bridging traditional navigation datasets into a standardized benchmark, named VLN-MME. We simplify the evaluation with a highly modular and accessible design. This flexibility streamlines experiments, enabling structured comparisons and component-level ablations across diverse MLLM architectures, agent designs, and navigation tasks. Crucially, enabled by our framework, we observe that enhancing our baseline agent with Chain-of-Thought (CoT) reasoning and self-reflection leads to an unexpected performance decrease. This suggests MLLMs exhibit poor context awareness in embodied navigation tasks; although they can follow instructions and structure their output, their 3D spatial reasoning fidelity is low. VLN-MME lays the groundwork for systematic evaluation of general-purpose MLLMs in embodied navigation settings and reveals limitations in their sequential decision-making capabilities. We believe these findings offer crucial guidance for MLLM post-training as embodied agents.
VLN-MME : Diagnosing MLLMs as Language-guided
@@
Xunyi Zhao, Gengze Zhou, Qi Wu Adelaide University, Australian Institute of Machine Learning {xunyi.zhao, gengze.zhou, qi.wu01}@adelaide.edu.au
@@ 1 Introduction
The rapid advancement of Multimodal Large Language Models (MLLMs) has catalyzed a shift from static vision-language tasks toward the development of embodied agents capable of dynamic, interactive decision-making. Central to this transition is the evaluation of 3D spatial reasoning. However, prior benchmarks have largely focused on static spatial Question Answering or simple navigation/manipulation tasks within low-fidelity, synthetic environments (yang2025thinking; daxberger2025mm; yang2025embodiedbench; cheng2025embodiedeval). Crucially, these settings lack substantial changes in 3D environmental states and do not require the maintenance of long-term spatial memory. In contrast, Vision-and-Language Navigation (VLN) (anderson2018r2r) emerges as a far more challenging paradigm for evaluating these abilities. Successfully navigating a 3D environment requires more than simple pattern recognition; it fundamentally tests an agent’s spatial understanding, its ability to maintain awareness, foresee the consequences of its actions across dynamic viewpoint changes, and the use of long-term memory to ground extended plans. Furthermore, when navigation involves multi-round dialogue, it probes the model’s capacity for contextual reasoning. However, despite VLN’s potential as a comprehensive benchmark for these core agentic skills, progress in systematically evaluating MLLMs is constrained by the limitations of existing evaluation pipelines.
@@
More critically, while recent evaluation suites like NavBench (qiao2025navbench) have attempted to standardize this process and successfully unify the evaluation of different tasks and models, however, they typically restrict to a single agent design. This limitation is substantial: by not systematically varying agent designs, it becomes impossible to decouple the intrinsic capabilities of the MLLM from the efficacy of specific prompting or planning strategies. As a result, the community lacks a fine-grained understanding of whether failures stem from the model’s reasoning limitations or suboptimal agent engineering.
In response to these challenges, we propose the Vision Language Navigation Multi-Model Evaluation (VLN-MME), a novel framework designed to address these limitations. Our approach features a modular and simulator-free architecture that preserves navigational semantics while eliminating the setup complexity and computational overhead of traditional pipelines. Unlike prior benchmarks, VLN-MME is designed to evaluate the triad of Model, Task, and Agent. Crucially, we move beyond high-level success metrics to provide a detailed error analysis, dissecting agent performance to diagnose core proficiencies in instruction following, spatial understanding, and historical sequential reasoning.
@@ 2 Related Works
Comprehensive benchmarks have emerged to test a wide spectrum of MLLM abilities (chaoyou2023mme; liu2024mmbench; li2024mvbench; yue2024mmmu; yu2024mm; lu2023mathvista; fei2025path), ranging from basic perception to complex cognition. Within this landscape, specific efforts have focused on assessing 3D spatial reasoning (yang2025thinking; daxberger2025mm; xu2025spatialbench; liao2024reasoning; li2024behavior). However, the majority of these benchmarks rely on static QA formats, where the model provides a single response to a fixed visual input, but not for continuous state tracking.
To assess sequential reasoning, several benchmarks focus on long-horizon tasks; however, these are largely restricted to digital domains like web browsing and application usage (deng2023mind2web; trivedi2024appworld; tao2025mmsearch; wang2025odysseybench). More recent embodied benchmarks attempt to bridge this gap but face distinct limitations. Some works, such as 3DMEM-Bench (hu20253dllm), focus primarily on high-level planning. While effective for evaluating abstract reasoning, these approaches often overlook the fine-grained environmental interactions required for realistic agent execution. Conversely, benchmarks like (yang2025embodiedbench; cheng2025embodiedeval) operate within low-fidelity, synthetic environments. Although they incorporate manipulation tasks, these settings generally lack photorealistic visual complexity and do not involve the massive 3D environmental state changes inherent to large-scale navigation. Consequently, these tasks rarely demand the long-term spatial memory or the rigorous sequential reasoning required to operate in dynamic, photorealistic spaces.
@@
The integration of MLLMs into robotics has inspired new paradigms for VLN. Early efforts leveraged LLMs as copilots to guide specialist agents (qiao2023march), while recent works employ off-the-shelf MLLMs as zero-shot agents through elaborate prompting (zhou2024navgpt; long2023discuss; chen2024mapgpt; dong2025se). Concurrently, other studies have explored fine-tuning MLLMs directly on navigation data (zhou2025navgpt2; lin2024navcot; pan2023langnav; zheng2023towards) or adapting pre-trained video models (zhang2024navid; zhang2024uni; cheng2024navila).
Despite this progress, evaluation remains fragmented and costly due to reliance on simulators. While recent efforts like SAME (zhou2025same) and NavBench (qiao2025navbench) attempt to standardize task evaluation, they face distinct limitations: the former does not specifically target MLLM agents, while the latter is restricted to fixed agent design. Consequently, existing frameworks often focus on aggregate metrics, lacking the fine-grained diagnostics required to understand specific failure modes. We bridge this gap with a unified, simulator-free framework that jointly evaluates MLLMs, agent designs, and diverse tasks.
@@ 3 Method
3.1 A Modular Framework for VLN Evaluation
@@ 3.2 Dataset Construction for Efficient Evaluation
To address the computational challenges of evaluating large models on existing, large-scale VLN datasets and to facilitate rapid experimentation, we constructed a curated benchmark for efficient yet representative evaluation. Follow the broader definition of VLN (zheng2023towards; zhou2025same), our benchmark is composed of samples carefully drawn from the validation unseen splits of three main datasets: R2R (anderson2018r2r), REVERIE (qi2020reverie), and ObjectNav (batra2020objectnav). Detailed descriptions of these datasets and the rationale behind their selection are provided in the Appendix A.
@@ 4.1 Settings
TL NE SR OSR SPL nDTW SDTW CLS TL NE SR OSR SPL nDTW SDTW CLS TL NE SR OSR SPL nDTW SDTW CLS
@@
GPT-5 9.12 6.28 38.50 57.50 29.23 37.96 22.02 34.11 9.37 8.03 30.00 49.50 20.76 27.11 14.88 21.05 9.63 4.41 48.00 79.00 23.84 26.52 16.33 20.13
Gemini-2.5 Pro 8.94 6.17 41.00 60.50 32.67 38.07 22.43 34.62 9.21 7.89 33.50 52.00 24.38 27.47 15.13 21.38 9.48 4.34 51.50 79.00 27.19 26.78 16.54 20.43
InternVL3-2B 9.89 8.56 13.50 27.00 5.46 21.25 5.75 21.59 10.13 10.18 7.33 16.33 2.50 15.30 2.97 17.17 10.35 6.27 21.50 40.50 3.57 13.67 2.55 15.57
InternVL3-8B 11.74 7.55 28.00 50.50 12.61 25.28 13.38 26.22 11.90 9.25 20.00 30.67 7.18 17.32 8.18 18.84 11.55 4.63 39.00 56.00 7.69 14.51 5.09 17.36
LLaVA-OV-7B 8.04 8.40 11.50 20.50 4.94 26.34 5.70 25.53 9.85 9.35 14.67 20.00 5.19 19.27 6.09 18.39 9.52 5.93 27.50 41.00 4.51 16.54 5.11 14.71
Qwen2.5-VL-7B 8.54 6.99 27.50 44.00 17.11 35.97 18.88 34.85 8.94 8.55 18.67 27.33 9.00 23.79 9.88 23.55 9.07 4.65 37.50 56.50 13.18 21.83 10.63 23.52
@@
GPT-5 9.47 6.42 36.00 54.50 27.31 39.17 18.89 32.54 9.71 8.24 28.50 46.00 20.42 32.53 12.04 26.82 9.92 4.58 46.50 70.00 23.67 31.06 14.03 25.07
Gemini-2.5 Pro 9.31 6.29 32.50 50.00 23.88 39.64 19.12 32.86 9.56 8.11 24.00 41.50 16.93 32.83 12.22 27.14 9.76 4.49 42.00 65.50 19.27 31.38 14.24 25.31
InternVL3-2B 6.21 8.84 8.00 14.50 4.47 26.15 4.28 27.45 4.98 9.87 5.33 9.00 3.24 23.39 3.11 25.09 6.78 5.56 25.00 41.00 6.66 25.51 5.34 23.98
InternVL3-8B 9.07 7.56 19.00 35.50 10.95 29.22 10.90 29.59 7.96 9.39 15.33 22.00 9.31 24.93 9.14 26.40 6.43 5.31 34.50 43.50 12.67 27.22 10.30 27.35
LLaVA-OV-7B 7.97 8.85 12.50 22.50 5.41 23.20 5.60 24.43 8.43 9.47 14.00 20.33 5.90 21.09 7.13 21.38 9.69 5.66 33.50 47.50 7.31 15.90 6.24 17.92
Qwen2.5-VL-7B 9.04 7.97 21.00 37.50 11.41 30.23 11.56 31.36 8.29 9.85 15.67 24.33 8.10 24.02 8.82 25.59 6.86 5.13 33.00 44.50 13.25 28.63 12.54 28.32
@@
GPT-5 9.61 6.37 33.00 50.50 24.18 40.04 19.53 33.14 9.82 8.13 25.00 43.00 17.34 33.03 12.51 27.08 10.03 4.47 43.00 68.00 20.62 30.02 15.06 24.09
Gemini-2.5 Pro 9.44 6.23 37.50 55.00 28.73 40.39 19.71 33.47 9.67 7.98 29.00 47.50 21.68 33.34 12.72 27.33 9.88 4.38 47.50 72.50 25.14 30.27 15.21 24.34
InternVL3-2B 6.50 8.94 8.00 16.00 5.20 29.91 5.00 30.13 7.20 9.41 8.50 15.00 4.80 25.49 4.60 24.93 6.43 5.56 28.00 43.00 7.83 26.53 6.14 25.28
InternVL3-8B 4.54 7.99 12.00 19.00 9.53 34.25 8.89 34.44 6.82 10.20 11.00 17.33 5.97 23.58 6.04 27.79 8.35 5.02 32.50 51.00 9.12 18.79 6.68 22.14
LLaVA-OV-7B 2.81 8.01 10.50 11.00 9.44 38.17 8.43 38.39 5.15 9.34 9.33 14.67 5.82 28.04 5.91 30.22 9.35 5.58 34.00 47.50 7.90 15.48 6.11 17.05
Qwen2.5-VL-7B 6.93 7.17 24.00 32.50 14.95 36.51 16.44 33.59 6.96 8.76 12.00 18.00 7.97 26.88 7.97 25.78 7.55 5.06 35.50 50.50 14.67 23.33 11.04 25.23
@@
GPT-5 9.52 6.31 38.50 56.50 29.81 39.56 20.08 33.02 9.77 8.18 30.00 48.50 22.17 32.04 13.07 26.03 9.97 4.36 48.50 73.50 25.92 29.08 15.51 23.06
Gemini-2.5 Pro 9.36 6.19 34.00 52.00 25.74 39.83 20.24 33.27 9.61 8.04 25.50 43.50 18.36 32.32 13.21 26.24 9.81 4.27 43.50 68.00 21.28 29.33 15.66 23.22
InternVL3-2B 7.15 9.24 4.50 15.00 1.70 22.45 2.05 23.47 7.30 9.78 9.33 15.00 4.63 22.33 5.30 24.07 6.94 6.25 24.50 37.50 7.26 21.29 6.11 21.47
InternVL3-8B 7.22 7.47 22.00 32.50 15.33 36.62 16.18 35.98 8.95 9.07 17.33 28.67 10.07 24.11 10.24 25.78 9.18 5.30 32.50 51.50 8.14 18.59 5.99 21.45
LLaVA-OV-7B 7.61 8.48 10.00 22.00 5.83 28.01 6.31 26.32 8.44 8.68 14.00 22.00 6.78 24.73 8.37 22.60 8.55 5.66 28.50 44.00 7.25 21.26 6.74 19.69
Qwen2.5-VL-7B 7.82 7.53 25.50 38.50 17.68 34.86 17.65 34.80 7.19 9.48 11.67 18.00 7.89 27.07 7.81 28.22 7.60 5.39 36.00 47.00 13.67 26.52 11.49 26.98
@@
GPT-5 9.32 6.21 34.00 52.50 25.83 39.03 21.08 33.05 9.53 7.94 26.00 45.00 18.29 29.07 13.02 23.04 9.73 4.33 44.00 74.00 20.91 28.06 16.08 22.01
Gemini-2.5 Pro 9.17 6.09 39.50 57.50 30.72 39.34 21.22 33.23 9.38 7.81 31.50 50.00 23.16 29.26 13.17 23.21 9.58 4.26 49.50 79.50 26.24 28.24 16.19 22.13
InternVL3-2B 9.84 8.61 11.00 22.50 3.71 20.18 3.85 21.89 10.19 9.59 12.00 19.00 4.41 18.10 5.24 19.97 10.35 5.93 27.50 46.50 4.41 13.59 3.69 15.57
InternVL3-8B 6.78 7.70 18.00 32.00 12.46 34.34 13.06 33.78 7.26 9.16 13.67 22.33 7.87 26.63 8.04 27.62 5.95 5.26 31.50 44.50 11.61 28.03 8.98 27.39
LLaVA-OV-7B 4.97 8.44 8.50 15.50 5.59 31.70 5.29 32.62 8.58 8.96 14.67 22.00 6.48 23.33 7.33 22.39 7.92 5.86 22.50 35.50 4.28 20.66 4.63 18.34
Qwen2.5-VL-7B 8.16 7.13 26.00 38.00 17.31 34.31 17.39 33.37 10.52 8.53 21.67 32.33 8.96 21.27 10.85 20.78 9.77 4.82 36.50 52.50 11.05 20.06 9.06 22.13
@@
GPT-5 9.62 6.51 32.50 50.00 27.14 40.09 17.04 32.08 9.83 8.31 24.50 42.50 19.68 34.05 10.06 28.03 9.93 4.67 44.50 68.50 23.21 32.08 13.01 26.02
Gemini-2.5 Pro 9.46 6.38 31.00 48.50 23.37 40.36 17.17 32.24 9.68 8.19 23.00 40.00 15.84 34.23 10.14 28.22 9.78 4.59 39.50 62.50 18.76 32.27 13.16 26.21
InternVL3-2B 5.09 8.99 6.50 14.00 4.12 26.43 3.94 27.62 5.22 9.88 4.00 8.00 1.58 23.01 1.77 24.48 6.66 6.31 20.00 35.50 4.76 21.14 4.69 21.05
InternVL3-8B 5.96 8.79 12.00 22.50 9.66 31.15 8.64 33.11 5.82 9.07 13.33 17.33 8.77 30.01 9.09 30.77 6.55 5.34 34.00 46.50 12.33 26.74 9.38 27.22
LLaVA-OV-7B 8.34 8.96 8.50 13.50 2.83 19.61 2.84 19.10 7.82 9.63 8.67 13.00 3.38 20.22 4.10 21.21 7.64 6.43 16.00 31.00 4.14 19.11 3.21 18.63
Qwen2.5-VL-7B 7.45 7.78 17.00 28.50 10.47 30.38 10.86 29.67 6.48 8.49 16.33 21.00 10.61 30.81 11.11 31.34 9.51 4.78 32.00 54.00 9.95 22.14 8.65 23.29

### v2

# Visual Navigation agents

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities across a wide range of vision-language tasks. However, their performance as embodied agents, which requires multi-round interaction with spatial reasoning and sequential action prediction, needs further exploration. Our work investigates this potential in the context of Vision-and-Language Navigation (VLN) by introducing a unified and extensible simulation-free evaluation framework to probe MLLMs as zero-shot agents, named VLN-MME. Simplifying the evaluation with a highly modular and accessible design streamlines experiments, enabling structured comparisons and component-level ablations across diverse MLLM architectures, agent designs, and navigation tasks. Crucially, enabled by VLN-MME, we observe that enhancing prevalent agents with Chain-of-Thought (CoT) reasoning and self-reflection leads to an unexpected performance decrease. This suggests MLLMs exhibit poor context awareness in embodied navigation tasks; although they can follow instructions and structure their output, their 3D spatial reasoning fidelity is low. Furthermore, we demonstrate that agent performance could be largely improved with simple failure cases in context learning. VLN-MME lays the groundwork for systematic evaluation of general-purpose MLLMs in embodied navigation settings and reveals limitations in their sequential decision-making capabilities. We believe these findings offer crucial guidance for MLLM post-training as embodied agents.
0.54902 0.08627 0.51765V0.59216 0.17255 0.52941L0.63922 0.25882 0.5451N0.68627 0.34118 0.55686-0.73333 0.42745 0.57255M0.77647 0.51373 0.58431M0.82353 0.6 0.6E\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: : Diagnosing MLLMs as Language-guided
@@
Xunyi Zhaothanks: Equal Contribution, $\dagger$ Project lead, Gengze Zhou∗†, Qi Wu Australian Institute for Machine Learning, Adelaide University {xunyi.zhao, gengze.zhou, qi.wu01}@adelaide.edu.au https://github.com/billzhao1030/VLN-MME
@@ 1 Introduction
The rapid advancement of Multimodal Large Language Models (MLLMs) has catalyzed a shift from static vision-language tasks toward the development of embodied agents capable of dynamic, interactive decision-making. Central to this transition is the evaluation of 3D spatial reasoning. However, prior benchmarks have largely focused on static spatial Question Answering or simple navigation or manipulation tasks within low-fidelity, synthetic environments (yang2025thinking; daxberger2025mm; yang2025embodiedbench; cheng2025embodiedeval). Crucially, these settings lack substantial changes in 3D environmental states and do not require the maintenance of long-term spatial memory. In contrast, Vision-and-Language Navigation (VLN) (anderson2018r2r) emerges as a far more challenging paradigm for evaluating these abilities. Successfully navigating a 3D environment requires more than simple pattern recognition; it fundamentally tests an agent’s spatial understanding, its ability to maintain awareness, foresee the consequences of its actions across dynamic viewpoint changes, and the use of long-term memory to ground extended plans. Furthermore, when navigation involves multi-round interaction, it probes the model’s capacity for contextual reasoning. However, despite VLN’s potential as a comprehensive benchmark for these core agentic skills, progress in systematically evaluating MLLMs is constrained by the limitations of existing evaluation pipelines.
@@
More critically, recent evaluation suites like NavBench (qiao2025navbench) have attempted to standardize this process and successfully unify the evaluation of different tasks and models. However, it is restricted to a single agent design and not systematically varying agent designs; it becomes impossible to decouple the intrinsic capabilities of the MLLM from the efficacy of specific prompting or planning strategies. As a result, the community lacks a fine-grained understanding of whether failures stem from the model’s reasoning limitations or suboptimal agent engineering.
In response to these challenges, we propose the Vision Language Navigation Multi-Model Evaluation (VLN-MME), a novel framework designed to address these limitations. Our approach features a modular and simulator-free architecture that preserves navigational semantics while eliminating the setup complexity and computational overhead of traditional pipelines. Crucially, we move beyond high-level success metrics to provide a detailed error analysis, dissecting agent performance to diagnose core proficiencies in instruction following, spatial understanding, and historical reasoning.
@@ 2 Related Works
Comprehensive benchmarks have emerged to test a wide spectrum of MLLM abilities (chaoyou2023mme; liu2024mmbench; li2024mvbench; yue2024mmmu; yu2024mm; lu2023mathvista; fei2025path), ranging from basic perception to complex cognition. Within this landscape, specific efforts have focused on assessing 3D spatial reasoning (yang2025thinking; daxberger2025mm; xu2025spatialbench; liao2024reasoning; li2024behavior). However, the majority of these benchmarks rely on static QA formats, where the model provides a single response to a fixed visual input, but not for continuous state tracking. To assess sequential reasoning, several benchmarks focus on long-horizon tasks; however, these are largely restricted to digital domains like web browsing and application usage (deng2023mind2web; trivedi2024appworld; tao2025mmsearch; wang2025odysseybench). More recent embodied benchmarks attempt to bridge this gap but face distinct limitations. Some works, such as 3DMEM-Bench (hu20253dllm), focus primarily on high-level planning. While effective for evaluating abstract reasoning, these approaches often overlook the fine-grained environmental interactions required for realistic agent execution. Conversely, benchmarks like (yang2025embodiedbench; cheng2025embodiedeval) operate within low-fidelity, synthetic environments. Although they incorporate manipulation tasks, these settings generally lack photorealistic visual complexity and do not involve the massive 3D environmental state changes inherent to large-scale navigation. Consequently, these tasks rarely demand the long-term spatial memory or the rigorous sequential reasoning required to operate in dynamic, photorealistic spaces.
@@
The integration of MLLMs into robotics has inspired new paradigms for VLN. Early efforts leveraged LLMs as copilots to guide specialist agents (qiao2023march), while recent works employ off-the-shelf MLLMs as zero-shot agents through elaborate prompting (zhou2024navgpt; long2023discuss; chen2024mapgpt; dong2025se). Concurrently, other studies have explored fine-tuning MLLMs directly on navigation data (zhou2025navgpt2; lin2024navcot; pan2023langnav; zheng2023towards) or adapting pre-trained video models (zhang2024navid; zhang2024uni; cheng2024navila). Despite this progress, evaluation remains fragmented and costly due to reliance on simulators. While recent efforts like SAME (zhou2025same) and NavBench (qiao2025navbench) attempt to standardize task evaluation, they face distinct limitations: the former does not specifically target MLLM agents, while the latter is restricted to fixed agent design. Consequently, existing frameworks often focus on aggregate metrics, lacking the fine-grained diagnostics required to understand specific failure modes. We bridge this gap with a unified, simulator-free framework that jointly evaluates MLLMs, agent designs, and diverse tasks.
@@ 3 Method
3.1 Modular VLN Evaluation Framework
@@ 3.2 Dataset Construction for Efficient Evaluation
To address the computational challenges of evaluating large models on existing, large-scale VLN datasets and to facilitate rapid experimentation, we constructed a curated benchmark for efficient yet representative evaluation. Following the broader definition of VLN (zheng2023towards; zhou2025same), our benchmark is composed of samples carefully drawn from the validation unseen splits of three main datasets: R2R (anderson2018r2r), REVERIE (qi2020reverie), and ObjectNav (batra2020objectnav). Detailed descriptions of these datasets and the rationale behind their selection are provided in the Appendix A.
@@ 4.1 Settings
TL NE nDTW SDTW CLS OSR SR SPL TL NE nDTW SDTW CLS OSR SR SPL TL NE nDTW SDTW CLS OSR SR SPL
@@
GPT-5 9.12 6.28 37.96 22.02 34.11 57.50 38.50 29.23 9.37 8.03 27.11 14.88 21.05 49.50 30.00 20.76 9.63 4.41 26.52 16.33 20.13 79.00 48.00 23.84
Gemini-2.5 Pro 8.94 6.17 38.07 22.43 34.62 60.50 41.00 32.67 9.21 7.89 27.47 15.13 21.38 52.00 33.50 24.38 9.48 4.34 26.78 16.54 20.43 79.00 51.50 27.19
InternVL3-2B 9.89 8.56 21.25 5.75 21.59 27.00 13.50 5.46 10.13 10.18 15.30 2.97 17.17 16.33 7.33 2.50 10.35 6.27 13.67 2.55 15.57 40.50 21.50 3.57
InternVL3-8B 11.74 7.55 25.28 13.38 26.22 50.50 28.00 12.61 11.90 9.25 17.32 8.18 18.84 30.67 20.00 7.18 11.55 4.63 14.51 5.09 17.36 56.00 39.00 7.69
LLaVA-OV-7B 8.04 8.40 26.34 5.70 25.53 20.50 11.50 4.94 9.85 9.35 19.27 6.09 18.39 20.00 14.67 5.19 9.52 5.93 16.54 5.11 14.71 41.00 27.50 4.51
Qwen2.5-VL-7B 8.54 6.99 35.97 18.88 34.85 44.00 27.50 17.11 8.94 8.55 23.79 9.88 23.55 27.33 18.67 9.00 9.07 4.65 21.83 10.63 23.52 56.50 37.50 13.18
@@
GPT-5 9.47 6.42 39.17 18.89 32.54 54.50 36.00 27.31 9.71 8.24 32.53 12.04 26.82 46.00 28.50 20.42 9.92 4.58 31.06 14.03 25.07 70.00 46.50 23.67
Gemini-2.5 Pro 9.31 6.29 39.64 19.12 32.86 50.00 32.50 23.88 9.56 8.11 32.83 12.22 27.14 41.50 24.00 16.93 9.76 4.49 31.38 14.24 25.31 65.50 42.00 19.27
InternVL3-2B 6.21 8.84 26.15 4.28 27.45 14.50 8.00 4.47 4.98 9.87 23.39 3.11 25.09 9.00 5.33 3.24 6.78 5.56 25.51 5.34 23.98 41.00 25.00 6.66
InternVL3-8B 9.07 7.56 29.22 10.90 29.59 35.50 19.00 10.95 7.96 9.39 24.93 9.14 26.40 22.00 15.33 9.31 6.43 5.31 27.22 10.30 27.35 43.50 34.50 12.67
LLaVA-OV-7B 7.97 8.85 23.20 5.60 24.43 22.50 12.50 5.41 8.43 9.47 21.09 7.13 21.38 20.33 14.00 5.90 9.69 5.66 15.90 6.24 17.92 47.50 33.50 7.31
Qwen2.5-VL-7B 9.04 7.97 30.23 11.56 31.36 37.50 21.00 11.41 8.29 9.85 24.02 8.82 25.59 24.33 15.67 8.10 6.86 5.13 28.63 12.54 28.32 44.50 33.00 13.25
@@
GPT-5 9.61 6.37 40.04 19.53 33.14 50.50 33.00 24.18 9.82 8.13 33.03 12.51 27.08 43.00 25.00 17.34 10.03 4.47 30.02 15.06 24.09 68.00 43.00 20.62
Gemini-2.5 Pro 9.44 6.23 40.39 19.71 33.47 55.00 37.50 28.73 9.67 7.98 33.34 12.72 27.33 47.50 29.00 21.68 9.88 4.38 30.27 15.21 24.34 72.50 47.50 25.14
InternVL3-2B 6.50 8.94 29.91 5.00 30.13 16.00 8.00 5.20 7.20 9.41 25.49 4.60 24.93 15.00 8.50 4.80 6.43 5.56 26.53 6.14 25.28 43.00 28.00 7.83
InternVL3-8B 4.54 7.99 34.25 8.89 34.44 19.00 12.00 9.53 6.82 10.20 23.58 6.04 27.79 17.33 11.00 5.97 8.35 5.02 18.79 6.68 22.14 51.00 32.50 9.12
LLaVA-OV-7B 2.81 8.01 38.17 8.43 38.39 11.00 10.50 9.44 5.15 9.34 28.04 5.91 30.22 14.67 9.33 5.82 9.35 5.58 15.48 6.11 17.05 47.50 34.00 7.90
Qwen2.5-VL-7B 6.93 7.17 36.51 16.44 33.59 32.50 24.00 14.95 6.96 8.76 26.88 7.97 25.78 18.00 12.00 7.97 7.55 5.06 23.33 11.04 25.23 50.50 35.50 14.67
@@
GPT-5 9.52 6.31 39.56 20.08 33.02 56.50 38.50 29.81 9.77 8.18 32.04 13.07 26.03 48.50 30.00 22.17 9.97 4.36 29.08 15.51 23.06 73.50 48.50 25.92
Gemini-2.5 Pro 9.36 6.19 39.83 20.24 33.27 52.00 34.00 25.74 9.61 8.04 32.32 13.21 26.24 43.50 25.50 18.36 9.81 4.27 29.33 15.66 23.22 68.00 43.50 21.28
InternVL3-2B 7.15 9.24 22.45 2.05 23.47 15.00 4.50 1.70 7.30 9.78 22.33 5.30 24.07 15.00 9.33 4.63 6.94 6.25 21.29 6.11 21.47 37.50 24.50 7.26
InternVL3-8B 7.22 7.47 36.62 16.18 35.98 32.50 22.00 15.33 8.95 9.07 24.11 10.24 25.78 28.67 17.33 10.07 9.18 5.30 18.59 5.99 21.45 51.50 32.50 8.14
LLaVA-OV-7B 7.61 8.48 28.01 6.31 26.32 22.00 10.00 5.83 8.44 8.68 24.73 8.37 22.60 22.00 14.00 6.78 8.55 5.66 21.26 6.74 19.69 44.00 28.50 7.25
Qwen2.5-VL-7B 7.82 7.53 34.86 17.65 34.80 38.50 25.50 17.68 7.19 9.48 27.07 7.81 28.22 18.00 11.67 7.89 7.60 5.39 26.52 11.49 26.98 47.00 36.00 13.67
@@
GPT-5 9.32 6.21 39.03 21.08 33.05 52.50 34.00 25.83 9.53 7.94 29.07 13.02 23.04 45.00 26.00 18.29 9.73 4.33 28.06 16.08 22.01 74.00 44.00 20.91
Gemini-2.5 Pro 9.17 6.09 39.34 21.22 33.23 57.50 39.50 30.72 9.38 7.81 29.26 13.17 23.21 50.00 31.50 23.16 9.58 4.26 28.24 16.19 22.13 79.50 49.50 26.24
InternVL3-2B 9.84 8.61 20.18 3.85 21.89 22.50 11.00 3.71 10.19 9.59 18.10 5.24 19.97 19.00 12.00 4.41 10.35 5.93 13.59 3.69 15.57 46.50 27.50 4.41
InternVL3-8B 6.78 7.70 34.34 13.06 33.78 32.00 18.00 12.46 7.26 9.16 26.63 8.04 27.62 22.33 13.67 7.87 5.95 5.26 28.03 8.98 27.39 44.50 31.50 11.61
LLaVA-OV-7B 4.97 8.44 31.70 5.29 32.62 15.50 8.50 5.59 8.58 8.96 23.33 7.33 22.39 22.00 14.67 6.48 7.92 5.86 20.66 4.63 18.34 35.50 22.50 4.28
Qwen2.5-VL-7B 8.16 7.13 34.31 17.39 33.37 38.00 26.00 17.31 10.52 8.53 21.27 10.85 20.78 32.33 21.67 8.96 9.77 4.82 20.06 9.06 22.13 52.50 36.50 11.05
@@
GPT-5 9.62 6.51 40.09 17.04 32.08 50.00 32.50 27.14 9.83 8.31 34.05 10.06 28.03 42.50 24.50 19.68 9.93 4.67 32.08 13.01 26.02 68.50 44.50 23.21
Gemini-2.5 Pro 9.46 6.38 40.36 17.17 32.24 48.50 31.00 23.37 9.68 8.19 34.23 10.14 28.22 40.00 23.00 15.84 9.78 4.59 32.27 13.16 26.21 62.50 39.50 18.76
InternVL3-2B 5.09 8.99 26.43 3.94 27.62 14.00 6.50 4.12 5.22 9.88 23.01 1.77 24.48 8.00 4.00 1.58 6.66 6.31 21.14 4.69 21.05 35.50 20.00 4.76
InternVL3-8B 5.96 8.79 31.15 8.64 33.11 22.50 12.00 9.66 5.82 9.07 30.01 9.09 30.77 17.33 13.33 8.77 6.55 5.34 26.74 9.38 27.22 46.50 34.00 12.33
LLaVA-OV-7B 8.34 8.96 19.61 2.84 19.10 13.50 8.50 2.83 7.82 9.63 20.22 4.10 21.21 13.00 8.67 3.38 7.64 6.43 19.11 3.21 18.63 31.00 16.00 4.14
Qwen2.5-VL-7B 7.45 7.78 30.38 10.86 29.67 28.50 17.00 10.47 6.48 8.49 30.81 11.11 31.34 21.00 16.33 10.61 9.51 4.78 22.14 8.65 23.29 54.00 32.00 9.95